# jmstate, a Flexible Python Package for Multi-State Joint Modeling

### v1

@@ 3.1 Graph Structure and Example
3.2 Individual Trajectories and the Markov Assumption
Each individual $i$ follows a true trajectory:
$\mathcal{T}_{i}^{*}=\left((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\ldots\right),$
where $T_{il}$ denotes the $l$-th transition time and $S_{il}\in V$ the corresponding state. Between transitions, the state $S_{i}(t)$ is assumed constant.
Only a finite portion of the trajectory is observed, up to a right-censoring time $C_{i}$:
$\big(\mathcal{T}_{i},C_{i}\big)=\Big(\big((T_{i0},S_{i0}),\dots,(T_{im_{i}},S_{im_{i}})\big),C_{i}\Big),$
where $m_{i}(t)=\max\{l:T_{il}\leq t\}$, and $m_{i}=m_{i}(C_{i})$ denotes the last observed transition index.
We assume that the process satisfies the (time-inhomogeneous) Markov property:
$\mathscr{L}\big((T_{i(l+1)},S_{i(l+1)})\mid\{(T_{il^{\prime}},S_{il^{\prime}})\}_{l^{\prime}\leq l}\big)=\mathscr{L}\big((T_{i(l+1)},S_{i(l+1)})\mid(T_{il},S_{il})\big).$ ($\ast$)
Remark 1.
The transition times are not restricted to be non-negative.
3.3 Transition Intensities and Survival Functions
Let $(k,k^{\prime})\in V^{2}$. The instantaneous risk of transition from state $k$ to state $k^{\prime}$ at time $t$, given entry at time $t_{0}$, is defined as:
$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})\coloneqq\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i(l+1)}\leq t+\delta,\ S_{i(l+1)}=k^{\prime}\mid T_{i(l+1)}>t,\ T_{il}=t_{0},\ S_{il}=k)}{\delta}.$
Summing over all admissible transitions from $k$, we get the total instantaneous risk of leaving state $k$:
$\sum_{s:\,(k,s)\in E}\lambda_{i}^{s\mid k}(t\mid t_{0})=\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i(l+1)}\leq t+\delta\mid T_{i(l+1)}>t,T_{il}=t_{0},S_{il}=k)}{\delta}.$
The cumulative intensity of the transition from $k$ to $k^{\prime}$ is defined by:
@@ 3.3 Transition Intensities and Survival Functions
which leads to the corresponding survival function:
$\mathbb{P}(T_{i(l+1)}>t\mid T_{il},S_{il})=\exp\left(-\sum_{s:\,(S_{il},s)\in E}\Lambda_{i}^{s\mid S_{il}}(t\mid T_{il})\right).$
Conditional transition probabilities are given by:
$\mathbb{P}(S_{i(l+1)}\mid T_{i(l+1)},T_{il},S_{il})=\frac{\lambda_{i}^{S_{i(l+1)}\mid S_{il}}(T_{i(l+1)}\mid T_{il})}{\sum_{s:\,(S_{il},s)\in E}\lambda_{i}^{s\mid S_{il}}(T_{i(l+1)}\mid T_{il})}.$
Combining both, the joint density is:
$p(T_{i(l+1)},S_{i(l+1)}\mid T_{il},S_{il})=\lambda_{i}^{S_{i(l+1)}\mid S_{il}}(T_{i(l+1)}\mid T_{il})\,\exp\left(-\sum_{s:\,(S_{il},s)\in E}\Lambda_{i}^{s\mid S_{il}}(T_{i(l+1)}\mid T_{il})\right).$
Furthermore, conditionning on the event $T_{i(j+1)}\geq t$ for some known $t$ can also easily be achieved by using the Chasles relation:
$\forall u\geq t,$
$\mathbb{P}(T_{i(l+1)}>u\mid T_{il},S_{il},T_{i(l+1)}>t)=\exp\left(-\sum_{s:\,(S_{il},s)\in E}\int_{t}^{u}\lambda_{i}^{s\mid S_{il}}(w\mid T_{il})\,dw\right).$
@@ 3.4 Multi-State Joint Model Specification
$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q),\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \lambda_{i}^{k^{\prime}\mid k}(t\mid X_{i},T_{im_{i}(t)},\mathcal{H}_{i}(t))=\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})\cdot\exp\big(\alpha^{k^{\prime}\mid k}g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})+\beta^{k^{\prime}\mid k}X_{i}\big),\end{cases}$
along with the assumptions 1, 2, and 5 as well as the following assumptions, analogous to assumptions 3 and 4:
Assumption 6.
Conditionally on $b_{i}$, the trajectories $\mathcal{T}_{i}^{*}$ are independent and satisfy the semi-Markov property $(\ast)$.
Assumption 7.
Conditionally on $b_{i}$, the biomarker process $Y_{ij}$ is independent of the transition process $\mathcal{T}_{i}^{*}$.
3.5 Clock Reset vs. Clock Forward
Two baseline hazard specifications are commonly used:
@@
The clock reset model resets the baseline hazard at each state entry and is appropriate when the transition risk depends on time spent in the state (Asanjarani et al. [2022]). The clock forward model assumes that risk accumulates with global time, ignoring the state entry time.
3.6 Initial State
We assume that the initial state $S_{i0}$ is observed. However, the framework could also incorporate a multinomial model for unobserved initial states (Yiu et al. [2018]).
This flexible framework enables inference on complex event histories, leveraging the full trajectory of biomarkers to refine predictions of transitions and survival [de Wreede et al., 2010].
3.7 Multi-State Simulation
The simulation of the transition process $\mathcal{T}_{i}^{*}$ conditionally on $b_{i}$, $X_{i}$, and $(T_{i0},S_{i0})$ can be achieved drawing on the semi-Markov property 6 by considering one transition at a time, according to Algorithm 1. The algorithme is similar to Gillespie’s algorithm [Gillespie, 2007], and can also include a survival condition $T_{i1}\geq t_{i}^{\text{surv}}$ that uses the Chasles relation LABEL:rel:chasles.
Algorithm 1 Simulation of trajectory $i$
2:
4:
6:  for each $s$ with $(S_{i(\ell_{i}-1)},s)\in E$ do
7:   Draw $T_{i\ell_{i}}^{s}$ such that
8:      $\log\mathbb{P}(T_{i\ell_{i}}^{s}>t)=\int_{T_{i(\ell_{i}-1)}\vee t_{i}^{\text{surv}}}^{t}\lambda_{i}^{s\mid S_{i(\ell_{i}-1)}}(w\mid\theta,X_{i},T_{i(\ell_{i}-1)},b_{i})\,dw,$
9:      $\forall t\geq T_{i(\ell_{i}-1)}\vee u.$
10:  end for
11:
13:  Append $(T_{i\ell_{i}},S_{i\ell_{i}})$ to $\mathcal{T}_{i}$.
14:  $\ell_{i}\leftarrow\ell_{i}+1$.
15:end while
16:
17:if $T_{i(\ell_{i}-1)}>C_{i}$ then
18:  Remove the last pair: $\mathcal{T}_{i}\leftarrow\mathcal{T}_{i}[:-1]$.
19:end if
[…]
4 Dynamic Prediction
We now focus on constructing random variables of interest for multi-state prediction for a new individual $i$, such that these random variables may be feasibly simulated with Algorithm 1.
Let $\tau_{i}$ be a stopping time adapted to the filtration $\mathcal{F}_{in}=\sigma(T_{i0},S_{i0},\dots,T_{in},S_{in})$, and let $\kappa_{i}=\{\inf n\in\mathbb{N}:\mathbb{P}(\tau_{i}=+\infty\mid\mathcal{F}_{in})=1\}$ be a second stopping time.
Assume that:
$\mathbb{P}(\min(\tau_{i},\kappa_{i})<+\infty)=1.$
Here, $\tau_{i}$ represents the time at which an event occurs for an individual $i$, while $\kappa_{i}$ is the time at which it becomes certain that the event will not occur. For instance, if the event of interest is patient remission, but the patient dies without remission, then $\tau_{i}=+\infty$, while $\kappa_{i}$ would be finite. As such, it is almost surely possible to simulate the trajectory of an individual $i$ up to the time of the event or the time of censoring.
Let $\xi$ be a function taking values in some space $\mathbb{Y}$:
$\xi:$ $(\mathbb{N}\cup\{+\infty\})^{2}\times(\mathbb{R}\times V)^{\mathbb{N}}\longrightarrow\mathbb{Y}$
$\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*}\longmapsto\xi(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=\xi\left(\tau_{i},\kappa_{i},(T_{il},S_{il})_{l\leq\min(\tau_{i},\kappa_{i})}\right)$
This function therefore only depends on the beginning of the trajectory and not on its entirety. Thus, during simulations, if $\tau_{i}$ is finite or almost surely infinite after a certain index, there is no need to simulate the entire trajectory. Applications include, for instance, the distributions of states at any time, transition times, return times, sojourn times, hitting times and many other quantities. This flexibility is illustrated by Examples 1 and 2.
Example 1 (State at time $t$).
Let $G=(V,E)$ be a directed acyclic graph, $t\in\mathbb{R}$ be a fixed time, and $\mathbb{Y}=V$ be the set of possible states. Let $\tau_{i}=\inf\{n\in\mathbb{N}:T_{in}\geq t\text{ or }S_{in}\text{ is absorbing}\}$. Clearly, $\tau_{i}\leq\operatorname{depth}(G)<+\infty$. Then, take $\xi_{t}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=S_{i\tau_{i}}$, which corresponds to the state of the individual $i$ at time $t$.
Example 2 (Hitting time).
Let $G=(V,E)$ be a directed acyclic graph, $\mathbb{A}\subset V$ a non-empty subset of states, and $\mathbb{Y}=\mathbb{R}\cup\{+\infty\}$. For any two non-empty sets $\mathbb{A},\,\mathbb{B}\subset V$, we note $\mathbb{A}\rightsquigarrow\mathbb{B}$ if there exists a path from $\mathbb{A}$ to $\mathbb{B}$ in $G$. Let $\tau_{i}=\inf\{n\in\mathbb{N}:S_{in}\in\mathbb{A}\}$, with the convention $\inf\emptyset=+\infty$. Then, $\kappa_{i}=\inf\{n\in\mathbb{N}:\{S_{in}\}\not\rightsquigarrow\mathbb{A}\}$ and $\min(\tau_{i},\kappa_{i})\leq\operatorname{depth}(G)<+\infty$. Finally, $\xi_{\mathbb{A}}(\tau_{i},\kappa_{i},\mathcal{T}_{i}^{*})=\mathds{1}_{\tau_{i}<+\infty}T_{i\tau_{i}}+\mathds{1}_{\tau_{i}=+\infty}\tau_{i}$ represents the hitting time for the set $\mathbb{A}$.
5 Likelihood and Statistical Inference
The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\text{full}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\text{marginal}}$, obtained by integrating them out:
$\mathcal{L}_{\text{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\text{full}}(\theta;;X,Y,\mathcal{T},C,b)\,db,$
where:
$\mathcal{L}_{\text{full}}(\theta;;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p(b_{i}\mid\theta)p(\mathcal{T}_{i},C_{i}\mid\theta,X_{i},b_{i})\prod_{j=1}^{m_{i}}p(Y_{ij}\mid\theta,X_{i},b_{i}).$
This expression is very similar to that obtained by Rizopoulos [2012] and is subdivided into three terms:
@@ 5 Likelihood and Statistical Inference
$p(Y_{ij}\mid\theta,X_{i},b_{i})=\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
@@
$p(\mathcal{T}_{i},C_{i}\mid\theta,X_{i},b_{i})={}$ $\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid\theta,X_{i},(T_{il},S_{il}),b_{i}\right)$ (1)
$\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{k^{\prime}\mid S_{im_{i}}}(C_{i}\mid\theta,X_{i},T_{im_{i}},b_{i})\right).$

### v2

@@ 3.2.1 Graph Structure
3.2.2 Individual Trajectories
Each individual $i$ follows a latent trajectory
$\mathcal{T}_{i}^{*}=\big((T_{i0},S_{i0}),(T_{i1},S_{i1}),(T_{i2},S_{i2}),\dots\big),$
where $T_{i\ell}$ denotes the $\ell$-th transition time and $S_{i\ell}\in V$ the corresponding state. The observed trajectory is right-censored at time $C_{i}$, so that only $\{(T_{i\ell},S_{i\ell}):T_{i\ell}\leq C_{i}\}$ is observed. Between transitions, the state $S_{i}(t)$ is piecewise constant. We note $m_{i}(t)\coloneq\sup\{\ell\geq 0:T_{i\ell}\leq t\}$ the index of the last observed transition before time $t$ and $m_{i}=m_{i}(C_{i})$.
The process is assumed to satisfy a (possibly time-inhomogeneous) semi-Markov property
$\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid\{(T_{i\ell^{\prime}},S_{i\ell^{\prime}})\}_{\ell^{\prime}\leq\ell}\big)=\mathscr{L}\big((T_{i,\ell+1},S_{i,\ell+1})\mid(T_{i\ell},S_{i\ell})\big),$ (1)
which is relaxed compared to a stricter Markov property by allowing sojourn-time dependence.
3.2.3 Transition Intensities and Survival Functions
For any admissible transition $(k,k^{\prime})\in E$, let $\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})$ denote the instantaneous risk of moving from $k$ to $k^{\prime}$ at time $t$, given entry time $t_{0}$
$\lambda_{i}^{k^{\prime}\mid k}(t\mid t_{0})=\lim_{\delta\to 0^{+}}\frac{\mathbb{P}(T_{i,\ell+1}\leq t+\delta,S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1}>t,T_{i\ell}=t_{0},S_{i\ell}=k)}{\delta}.$
The cumulative intensity and the corresponding survival function are
@@ 3.2.3 Transition Intensities and Survival Functions
$\mathbb{P}(T_{i,\ell+1}>t\mid T_{i\ell},S_{i\ell})=\exp\left[-\sum_{s:(S_{i\ell},s)\in E}\Lambda_{i}^{s|S_{i\ell}}(t\mid T_{i\ell})\right].$
Conditional transition probabilities are
$\mathbb{P}(S_{i,\ell+1}=k^{\prime}\mid T_{i,\ell+1},T_{i\ell},S_{i\ell})=\frac{\lambda_{i}^{k^{\prime}|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}{\sum_{s:(S_{i\ell},s)\in E}\lambda_{i}^{s|S_{i\ell}}(T_{i,\ell+1}\mid T_{i\ell})}.$
@@ 3.3 Multi-State Joint Model Specification
$\begin{cases}Y_{ij}=h(t_{ij},\psi_{i})+\epsilon_{ij},\quad\epsilon_{ij}\sim\mathcal{N}(0,R),\\ \psi_{i}=f(\gamma,X_{i},b_{i}),\quad b_{i}\sim\mathcal{N}(0,Q),\\ \lambda_{i}^{k^{\prime}\mid k}\left(t\mid X_{i},T_{im_{i}(t)},\mathcal{H}_{i}(t)\right)=\lambda_{0}^{k^{\prime}\mid k}(t\mid T_{im_{i}(t)})\exp\big(\alpha^{k^{\prime}\mid k}g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})+\beta^{k^{\prime}\mid k}X_{i}\big).\end{cases}$
Here, $h$ and $f$ define the nonlinear mixed-effects submodel, $g^{k^{\prime}\mid k}$ represents the link between biomarker dynamics and the transition intensity, and $\lambda_{0}^{k^{\prime}\mid k}$ denotes the baseline hazard associated with transition $(k\to k^{\prime})$. This structure extends the classical joint model to arbitrary directed graphs and to both Markovian and semi-Markov specifications. The coefficient $\alpha^{k^{\prime}\mid k}$ quantifies the effect of the longitudinal biomarker dynamics on the instantaneous risk of transitioning from state $k$ to state $k^{\prime}$. More concretely, for a one-unit increase in $g^{k^{\prime}\mid k}(t,X_{i},\psi_{i})$ (holding other covariates constant), the hazard ratio for transition $(k\to k^{\prime})$ is $\exp(\alpha^{k^{\prime}\mid k})$. Likewise, $\beta^{k^{\prime}\mid k}$ represents the effect of baseline (or time-varying) covariates $X_{i}$ on that same transition: a one-unit increase in a covariate $X_{ij}$ multiplies the hazard by $\exp(\beta^{k^{\prime}\mid k}_{j})$. Thus, $\alpha^{k^{\prime}\mid k}$ captures the degree to which the biomarker trajectory (via its latent parameter $\psi_{i}$ or function $g^{k^{\prime}\mid k}$) influences the transition risk, while $\beta^{k^{\prime}\mid k}$ captures the direct effect of covariates on that transition’s risk, beyond the biomarker pathway.
3.3.1 Model Variants
Two baseline hazard conventions are common:
@@ 3.3.1 Model Variants
The clock-reset form models risk as a function of time spent in the current state, whereas the clock-forward form measures risk with respect to global time since study entry.
3.3.2 Likelihood Formulation
To derive the marginal likelihood, we decompose the joint distribution of the longitudinal and multi-state processes under a set of standard conditional independence assumptions detailed in Table 3.3.2 below.
Assumptions for Likelihood Factorization A. Latent-level independence A1. Random effects $(b_{i})_{i}$ are mutually independent across individuals. B. Conditional independence within individuals B1. Longitudinal observations $(Y_{ij})_{ij}$ are mutually independent given $b_{i},X_{i}$. B2. Trajectories $(\mathcal{T}_{i}^{*})_{i}$ are mutually independent given $b_{i},X_{i}$. B3. Conditional on $b_{i},X_{i}$, the longitudinal and event processes are mutually independent. C. Censoring and process assumptions C1. Censoring times $(C_{i})_{i}$ are mutually independent and noninformative, i.e. $C_{i}\mathrel{\perp\!\!\!\perp}(Y_{i},\mathcal{T}_{i}^{*})\mid b_{i},X_{i}$. C2. Conditionally on $b_{i},X_{i}$, event trajectories satisfy the semi-Markov property 1: $p(\mathcal{T}_{i}^{*}\mid b_{i},X_{i})=\prod_{\ell\geq 0}p\left((T_{i,\ell+1},S_{i,\ell+1})\mid T_{i\ell},S_{i\ell},b_{i},X_{i}\right)$.
These grouped assumptions mirror those used in classical joint modeling [rizopoulos2012joint] and multi-state survival analysis [putter2007competing]: independence across subjects; the conditional independence structure linking the longitudinal and event submodels through shared random effects; noninformative censoring and (semi-)Markovian dynamics.
Let $\theta=(\gamma,Q,R,\alpha,\beta)$ denote the vector of model parameters. For subject $i$, we observe the longitudinal measurements $Y_{i}=(Y_{i1},\dots,Y_{in_{i}})$ and the event trajectory $\mathcal{T}_{i}=\{(T_{i\ell},S_{i\ell})_{0\leq\ell\leq m_{i}}\}$. The joint likelihood then factorizes as
$p(Y_{i},\mathcal{T}_{i}\mid X_{i},C_{i},\theta)=\int p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta)\,db_{i},$
where $b_{i}$ are individual random effects. This formulation generalizes the joint likelihoods of wulfsohn1997joint and ferrer2016joint to arbitrary multi-state event structures. Each part can then be explicitely expressed using Assumptions 3.3.2, yielding an expression very similar to that obtained by rizopoulos2012joint.
@@ 3.3.2 Likelihood Formulation
$p(Y_{i}\mid b_{i},X_{i},\theta)=\prod_{j=1}^{n_{i}}\frac{(2\pi)^{-d/2}}{\det(R)^{1/2}}\exp\left(-\tfrac{1}{2}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)^{T}R^{-1}\left(Y_{ij}-h(t_{ij},\psi_{i})\right)\right),$
@@
$p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)=$ $\prod_{l=0}^{m_{i}-1}p\left((T_{i(l+1)},S_{i(l+1)})\mid b_{i},X_{i},(T_{il},S_{il}),\theta\right)$ (2)
$\exp\left(-\sum_{s:\,(S_{im_{i}},s)\in E}\Lambda_{i}^{k^{\prime}\mid S_{im_{i}}}(C_{i}\mid b_{i},X_{i},T_{im_{i}},\theta)\right).$
The proof of the expression of the Semi-Markov likelihood (2) is provided in Appendix B.
The likelihood factorization above is valid and forms the basis for scalable inference procedures described in Section 4, leveraging the complete trajectory of biomarkers to refine predictions of transitions and survival [de2010mstate].
Moreover, here, we assume that the initial state $S_{i0}$ is observed. However, the framework could also incorporate a multinomial model for unobserved initial states [yiu2018clustered].
3.4 Model Selection and Information Criteria
To compare competing model specifications, we rely on information criteria based on the marginal likelihood as defined in Section 4. The Akaike Information Criterion (AIC) [akaike2003new] and the Bayesian Information Criterion (BIC) [schwarz1978estimating] are standard tools. While the AIC relies on an asymptotically unbiased estimator of the $\log$-likelihood, the BIC aims to identify the true model with great probability as the sample size grows.
The Akaike Information Criterion is computed as follows
$\mathrm{AIC}=-2\log\mathcal{L}_{\text{marginal}}(\hat{\theta};X,Y,\mathcal{T},C)+2k,$
where $k$ is the number of parameters, and $\hat{\theta}$ is a maximizer of the marginal likelihood.
The Bayesian Information Criterion is similarly defined but imposes a stronger penalty for model complexity, growing in $\log n$ where $n$ is the number of observations. Particularly in the context of mixed effects models, and even more so for multi-state joint models, the number of observations $n$ may vary from one definition to another, either the total number of repeated measurements or the number of individuals [delattre2014note]. To alleviate this problem, we recall the derivation of the BIC from the Laplace approximation [lebarbier2006introduction]. For a prior distribution $\pi$ on a set of models $\mathcal{M}$, the Laplace approximation approaches up to constant terms which could also be accounted to improve the approximation, the posterior probability given observed data $x$ defined in Section 4 by
@@ 3.4 Model Selection and Information Criteria
where $H_{\hat{\theta}}$ denotes the Hessian matrix of the marginal $\log$-likelihood evaluated at the Maximum Likelihood Estimator (MLE). Then, under standard regularity conditions that allow differentiation and integration to be interchanged twice, we have that
[…]
Thus, our BIC criterion can be written as
[…]
In both cases, one should aim at minimizing the chosen criterion.
4 Statistical Inference
The estimation of model parameters $\theta=(\gamma,Q,R,\alpha,\beta)$ relies on two likelihood formulations: the complete-data likelihood $\mathcal{L}_{\text{complete}}$, which includes latent variables, and the marginal likelihood $\mathcal{L}_{\text{marginal}}$, obtained by integrating them out
$\mathcal{L}_{\text{marginal}}(\theta;X,Y,\mathcal{T},C)=\int\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)\,db,$
where
$\mathcal{L}_{\text{complete}}(\theta;X,Y,\mathcal{T},C,b)=\prod_{i=1}^{n}p(Y_{i}\mid b_{i},X_{i},\theta)\,p(\mathcal{T}_{i}\mid b_{i},X_{i},C_{i},\theta)\,p(b_{i}\mid\theta).$